OpenStudy (anonymous):

square root of 27 +(plus) the square root of 75 and Why

OpenStudy (anonymous):

\[\sqrt{27}+\sqrt{75}\]Now what do we know about 27 and 75. Can you think of any factors?

OpenStudy (anonymous):

5?

OpenStudy (anonymous):

\[\sqrt{9 \times 3} + \sqrt{3 \times 25}\]

OpenStudy (anonymous):

which we can treat as \[ 3\sqrt{3} + 5\sqrt{3}\]

OpenStudy (anonymous):

oh! i get it. Thanks!

OpenStudy (anonymous):

Which becomes \[8\sqrt{3}\]

OpenStudy (anonymous):

great! No problem